OpenStudy (angela210793):

Some help plsss Point K has distance 15cm from the vertexes of a triangle ABC.BC = 8cm and m (BAC) = 30 degrees, Find the distance from K to the plan of the triangle ...

OpenStudy (anonymous):

which is the vertex ? A or B or C ?

OpenStudy (angela210793):

it has the same distance from all the 3 vertexes =15

OpenStudy (anonymous):

oh sry my bad

OpenStudy (angela210793):

it's ok...can u help me?

OpenStudy (anonymous):

what is meant by the plan of the triangle ?

OpenStudy (angela210793):

the plan where the triangle is....

OpenStudy (angela210793):

hmm..let me see if i can draw it

OpenStudy (anonymous):

well i already solved the triangle

OpenStudy (angela210793):

u should know that KO is perpendicular with the plan of triangle and i guess it has to do with a circle...but idk wht circle it is

OpenStudy (anonymous):

so i guess i can help if i know what kind of plan your problem means

OpenStudy (angela210793):

see my attach. maybe it'll help :)

OpenStudy (anonymous):

lmao thats a prism :)

OpenStudy (angela210793):

Hey...come on...ofc would be a prism if u joined the points..wht did ya think????? :P:P

OpenStudy (anonymous):

i thought a triangle :P

OpenStudy (angela210793):

Lol :D:D ...how abt now tht u got the pic...do u think u can solve it? Ps:I'm not 100% sure that O is in there

OpenStudy (anonymous):

idk honestly are u sure its talking about a prism ?

OpenStudy (angela210793):

:@@@@ Nooo!!!!!.... U've never solved this kind of problem in geometry? But's ok :) thnx 4 trying anyway :):)

OpenStudy (anonymous):

cause there is infinite values of regular distances to the plan that u showed

OpenStudy (angela210793):

Nooo....how infinite?

OpenStudy (anonymous):

unless it asked to find the perp. distance aka the shortest dist (and it didnt)

OpenStudy (angela210793):

Hmmm it asks to find the perpendicular...KO is perpendicular with the triangle's plan...so it asks tht...

OpenStudy (anonymous):

don'tgot the question, is that a pyramid or triangle??

OpenStudy (anonymous):

cant solve sry

OpenStudy (angela210793):

it's ok :):) I'll ask some1 else..thnx a loot 4 trying though :):):) A prism Korcan

OpenStudy (anonymous):

u say I am a prism??

OpenStudy (angela210793):

Huh? U asked if it was a triangle or a pyramid and i said it is a prism!!!! :P

OpenStudy (anonymous):

ok let me have a look ^^

OpenStudy (anonymous):

just use pisagor...

OpenStudy (anonymous):

And i can't draw it on to my notebook :P its hard to draw an triangle prish lol...

OpenStudy (angela210793):

u can't draw it????? :P:P C'mon..don't b mean..I've seen wht u can do remember???????? :P:P

OpenStudy (anonymous):

yes u saw but did u see any triangle prism on my draws?? :P

OpenStudy (angela210793):

Hahahhaha :P:P:P Well there's always a 1st time...so.....

OpenStudy (anonymous):

But its hard!!

OpenStudy (angela210793):

How did u draw it? Tht's not hard...just use imagination....

OpenStudy (anonymous):

wekk ş draw one but in tthat k was under the ABC triangle :P

OpenStudy (angela210793):

kk...it'll b the same :P:P:P

OpenStudy (anonymous):

:P Butt its really HARd to draw !!! :((

OpenStudy (angela210793):

Lool :P:P:P Ok...4get it :P:P:P

OpenStudy (anonymous):

ok ^^

OpenStudy (anonymous):

i am good at drawing^3D pics... but can't draw it with pencil ^^ betteri n comp

OpenStudy (angela210793):

:P:P:P

OpenStudy (anonymous):

hmmm some one can solve this yet ?

OpenStudy (anonymous):

i'm on it, working will giv u the answer :)

OpenStudy (anonymous):

hope so cause i couldnt solve it either

OpenStudy (anonymous):

hmmm dont u believe that since KA = KB = KC then OA = OB = OC ?

OpenStudy (angela210793):

Yea i got a friend he's a genius :):):) Is just tht he's not online 4 now :):)

OpenStudy (angela210793):

yea u see in the post above i told u tht it has to do with a circle....so u r thinking the same right?

OpenStudy (angela210793):

@brackett Thnx :)

OpenStudy (anonymous):

well yes and another thing

OpenStudy (anonymous):

for OB = OC = OB aint AC = AB ?

OpenStudy (anonymous):

basically consider a sphere of radius 15 , passing through the vertices of the triangle. u need to find the distance of the centre of the sphere from the plane of the triangle :)

OpenStudy (angela210793):

Yea..tht's wht i thought Brackett but still i can't solve it...how abt the angle=30 given? where to use it?

OpenStudy (dumbcow):

ok solution is KO = 12.689 Like bnb said since KA=KB=KC then OA=OB=OC This divides triangle ABC into 3 isosceles triangles where each isosceles triangle has 2 equal bottom angles, which are A,B,C split into 2 I labeled them accordingly: mA = x+y mB = y+z mC = x+z ex: the triangle BOC with 8 as its base has angles z Now the sum of these angles equal 180 2x+2y+2z = 180 x+y+z = 90 x+y = 30, given mA=30 30+z = 90 z=60 thus triangle BOC is equilateral and all sides are 8 OA=OB=OC = 8 Now look at right triangle KOA OA=8, KA=15 KO = sqrt(15^2 - 8^2) = 12.689

OpenStudy (dumbcow):

this took me awhile, geometry ugh .. :)

OpenStudy (anonymous):

dumbcow: that was stud :) easy proof i used vectors :)

OpenStudy (angela210793):

Woooooooooow!!! It took me all morning and i couldn't solve it at all...Thnx a looooot :):):) Thank u all 4 trying :):)

OpenStudy (dumbcow):

your welcome

OpenStudy (angela210793):

^_^

OpenStudy (anonymous):

good one

OpenStudy (angela210793):

I got 10 problems similar to this...but thnx God i know how to solve the other 9 :))